OpenStudy (anonymous):

The number of mulch you need to cover a planting area varies jointly with the area to be mulched a in square feet and the depth of the mulch d in feet. If you need 10 bags to mulch 120ft2 to a depth of 3in, how many bags do you need to mulch 200ft2 to a depth of 4in?

OpenStudy (anonymous):

Make sure you include the model.

OpenStudy (anonymous):

\[ 10n = 120ft^2\times 3in = 120(12in)^2 \times 3in \]Solve for \(n\). Then: \[ m\times n = 200ft^2 \times 4in = 200(12in)^2 \times 4in \]Solve for \(m\).

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Could you please help me with another question?

OpenStudy (anonymous):

Gotta go soon

OpenStudy (anonymous):

Okay. Thanks anyways.